Victory Bank reports an EM​ (equity multiplier) for 30 while Batovi Bank reports an EM equal to 18. The equity capital as a percentage of assets for Victory Bank is??? and Batovi Bank
Answer #1

Equity multiplier = Assets/ Equity

Victory Bank

equity capital as a percentage of assets = 1/EM = 1/30 = 0.0333 = 3.33%

Batovi Bank

equity capital as a percentage of assets= 1/EM = 1/18 = 0.0556 = 5.56%